Grade 11General Physics

A particle is thrown with a speed of 60 m/s at an angle 60 degree to the horizontal. When the particle makes an angle 30 degree with the horizontal in downward direction its speed at that instant is V . what is value of V.

Kalpit
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
velocity of particle is given by
vx = 60cos60 = 30 m/s
vy = 60sin60 – gt = 30\sqrt{3} – gt
SO v = 30\hat{i} + (30\sqrt{3} – gt)\hat{j}
and angle with horizontal is 30 in down ward direction
vcos30 = 30
So v = 60/\sqrt{3}
v = 20\sqrt{3} m/s
Hope it clears.